OpenStudy (anonymous):

Write a program that computes and prints the 1000th prime number, can somebody help me?

OpenStudy (anonymous):

Show us what you got so far and we can work from there. It's better to give it a try first rather than just read someone's solution.

OpenStudy (anonymous):

x=1 y=2 z=0 while z<10: if y<=x: if x%y==0: print x, x=x+2 y=2 z=z+1 else: y=y+1 else: x=x+2 z=z+1 y=2 i wrote this, try to run this and try to find the errors.thank you

OpenStudy (anonymous):

Here is what I want you to do: Just before Print x, I want you to add Print x, y, x%y. You should notice y isn't doing exactly what you want it to do (reaches x instead of x-1) When trying to debug a program always add in Print statements. Print the variables and try to see if they match up to what you expect them to be. If not, try to figure out why it happened. If you need some more help with this particular problem let me know.

OpenStudy (anonymous):

what pogramming language is this???

OpenStudy (anonymous):

python

OpenStudy (anonymous):

kNicola thank you i finally find the solution by using the "for loop' rather than while

OpenStudy (anonymous):

No problem. You could have also added an elif clause. elif y < x: y=y+1

OpenStudy (anonymous):

Come on now, without giving it any thought at all, you can just cycle through the integers checking whether they are prime or not. Keep track of how many primes you have found so far. When you get to the thousandth one or whatever it is, return it.